OpenStudy (anonymous):

A basket contains 11 pieces of fruit: 4 apples, 5 oranges, and 2 bananas. Jonas takes a piece of fruit at random from the basket, and then Beth takes a piece at random. What is the probability that Jonas will get an orange and Beth will get an apple?

OpenStudy (anonymous):

9/10?

OpenStudy (anonymous):

Total outcomes = 11 favorable outcomes = 5 Probability = Number of favorable outcomes / Total number of outcomes Probability that Jonas will get an orange = 5/11 Total outcomes = 10 (because 1 fruit is already taken by Jonas) favorable outcomes = 4 Probability Beth will get an apple = 4/10 = 2/5

OpenStudy (anonymous):

so 9/11?

OpenStudy (anonymous):

no , multiply both probability that Jonas will get an orange and Beth will get an apple = 5/11 * 2/5 = 2/11